OpenStudy (anonymous):

The supplement of an angle measures 15 more than four times its complement . Find the measure of the angle .

OpenStudy (nurali):

a = angle. x = complement. 4x + 15 = supplement. a + x = 90 a + 4x + 15 = 180 180 - 90 = 90 so a + 4x + 15 - a - x = 90 3x + 15 = 90 (a cancels out and 4x becomes 3x after subtration) 3x = 75 (subtracting 15 from both sides) x = 25 (dividing both sides by 3) if x = 25, then a = 65 because 65 + 25 = 90. if x = 25, then 4x + 15 = 115 + a = 180, so a = 65 again because 65 + 115 = 180. angle is 65 degrees.

Directrix (directrix):

Let x be the angle. --> Supplementary angles total 180. The supplement of x is 180 - x. --> Complementary angles total 90. The complement of x is 90 - x. -->

Directrix (directrix):

Let x be the angle. The supplement of x is 180 - x. The complement of x is 90 - x.

Directrix (directrix):

The supplement of an angle measures 15 more than four times its complement . --------------------------------- Translate above words to symbols. (supplement of an angle) is 15 (more than) (four times its complement) 180 - x = 15 + 4( 90 - x) ---> Solve this for x. @janhell123

OpenStudy (nurali):

if correct please click best response.